LOG-IN
Displaying report 1-1 of 1.
MIF (ITF Control)
Print this report.
yutaro.enomoto - 22:37, Tuesday 27 November 2018 (7150)Get code to link to this report
Switching off them all: Noise performace of ALS improved

[Miyoki, Nakano, Miyazaki, Enomoto]

We switched off all the air flow related things we can stop and checked the control signal spectrum of Gr PDHX. And fount it improved!

We switched off KOACH for PSL room, air temerature controllers for PSL, an air compressor between PR2/3, and FFU clean filters for PR-MCF area and BS area.
Here is the comparison of the control signals with various conditions.
Black trace corresponds to all-on condition, red one to all-off condition, and cyan one corresponds to FFU for the condition where PR-MCF area was on. Blue one corresponds to the error signal of IMC loop taken simaltaneously with cyan one.
Improvment above ~200Hz comes from switching off the KOACH, and improvement below 100Hz comes from switching off air temerature controllers for PSL (and partially from FFUs).
It turned out that operating one of the air temperature controllers does not increase the noise significantly.

And then, we closed ALS CARM loop and saw if we can keep the IR resonance or not.
Here shows the transmission of IR when the loop was closed. This became significantly stable.
The condition for the following figure: one air temperature controller was on and the air duct was half-opened.

Note:
-- From the PDHX control signal spectra, it apears that FFUs do not have much influence on ALS, but they do. By switching off FFUs, the IR tranmission was calmed down very much. This implies that fiber phase noise induced by the vibration of the cable rack limits the performanc of ALS.
-- How air temperature controllers induce noise on ALS is not so clear, but one possibility is through the vibration of the wall of PSL room, similarly to the above noise path.
-- We still see the frequency fluctuation in IR transmission, at several tens of Hz and ~0.2-0.5 Hz. Still do not know where they come from. Not enough gain in microseismic region? some scattering? just another fiber phase noise?

Images attached to this report
7150_20181127143636_alsswoff.png 7150_20181127143655_keepirpslairconditionerhalf1unit.png
Comments related to this report
kiwamu.izumi - 23:02, Tuesday 27 November 2018 (7151)

Fiber phase noise after all... Great job, gentlemen!

By the way, how long were you able to hold the IR within a resonance?

osamu.miyakawa - 23:21, Tuesday 27 November 2018 (7154)

Wow! Very goooood results!

shinji.miyoki - 07:29, Wednesday 28 November 2018 (7157)
Uchiyama-kun pointed out the low frequency noise might come from the bottom plate low frequency motion where the last mirror is fixed for the green beam injection. I think it could be.

If it has angular motion and the green beam axis is off-center for the 3km cavity, these coupling produces beam jitter noise which the only green beam senses for the low fequency range. If this might be the reason, the performance difference bw IR and GR for the low frequency range will be expected.
shinji.miyoki - 23:57, Tuesday 27 November 2018 (7156)
Akutsu-kun pointed out the designed performance of pylon is comparable with the drift that we observed.
yutaro.enomoto - 09:43, Wednesday 28 November 2018 (7160)

> By the way, how long were you able to hold the IR within a resonance?

Typically ~ 1 or 2 seconds without the transmission touching zero.